Reports 1-1 of 1 Clear search Modify search
Detector Operation (Operations Report)
Montanari - 16:00 Tuesday 22 January 2019 (44481) Print this report
Operator Report -Morning Shift Jan 22

This morning I found the ITF locked in LOW NOISE 1.

The first part of the shift was dedicated to the maintenance activities:

INF              Standard Cleaning Operations;
                    Oxygen deficiency sensors replacement in Terminal Buildings (external firm);
VACUUM    Standard Nytrogen Refill;
ENV             I.Fiori went in Laser Lab to install and relocated some accelerometers;
TCS             Chiller Check performed at 08:45UTC, the refill has been needed only for the NI one;
SUSP          Sa_PR cooling unit replacement.

After that, started the recovery and still in progress.

VPM
10:49UTC Group Earthquakes processes restarted after a crash
11:29UTC USGS process restarted after a crash

Search
×

Warning

×